# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v3

@@ Abstract
Real-time corridor-wide crash occurrence risk (COR) prediction is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to capture spatial heterogeneity in traffic and roadway conditions. To address these gaps, this study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator for capturing V–V and V–I near misses, and a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) UGEV model for estimating short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset were analyzed across 28 sites on Miami’s Biscayne Boulevard to extract extreme near-miss events. The model incorporates vehicle dynamics and roadway features as covariates, with partial pooling across segments and intersections to capture corridor-wide heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near misses. Predictive validation using ROC–AUC confirms strong accuracy (0.89 for V–V segments, 0.82 for intersections, 0.79 for V–I segments, and 0.75 for intersections). Random-parameter effects showed that relative speed, distance, and deceleration dominated V–V near misses at segments (only distance for V–I), while intersections were dominated by relative speed and distance for V–V, with no significant effects for V–I.. These findings demonstrate the value of a geometry-aware, spatially adaptive framework for proactive corridor safety management, supporting both real-time interventions and long-term Vision Zero goals.
@@ 1 Introduction
The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches. Urban corridors are particularly vulnerable, as even a single crash can disrupt mobility, trigger cascading congestion, and result in substantial economic losses. In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes exceeds \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite decades of research, corridor-wide crash risk prediction remains methodologically limited.
@@
Crash risk prediction has traditionally relied on retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011). While such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions. These drawbacks highlight the need for proactive approaches that anticipate rather than react to crash risk.
@@
To address these challenges, researchers have shifted toward proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among these, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, their assumptions of constant velocity and linear paths bias results in boundary-constrained settings. Recent two-dimensional (2D) extensions (Li et al., 2024; Anis et al., 2025b) incorporate position, heading, and acceleration, yet still rely on simplified circular vehicle approximations. This study addresses these limitations by introducing a geometry-aware 2D-TTC that accounts for vehicle dimensions, dynamics, and roadway boundaries to realistically capture both vehicle–vehicle (V–V) and vehicle–infrastructure (V–I) near misses.
@@
Building on this foundation, recent studies have integrated SSMs with extreme value theory (EVT) to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). EVT extrapolates from frequent near misses to rare crash-prone extremes, offering a statistically principled framework for proactive safety analysis. Yet most SSM–EVT applications remain limited, focusing narrowly on V–V interactions and excluding V–I conflicts such as lane edges, curbs, and barriers. This omission is critical because V–I events account for a large share of severe roadway departure crashes (Islam et al., 2021). Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable precise alignment of roadway boundaries with vehicle trajectories, creating new opportunities to model both V–V and V–I near misses with geometric fidelity.
@@
Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across study areas. More recent work has developed nonstationary GEV formulations, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and capture heterogeneity across traffic states. (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions further address data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). Despite these advances, most applications remain geographically narrow and functionally restricted to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have integrated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), explicit corridor-wide models that jointly integrate vehicle dynamics, roadway geometry, and directional heterogeneity remain rare.
To address this gap, this paper makes three contributions: (1) it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions, (2) it introduces a geometry-aware 2D-TTC indicator that captures extreme near-miss dynamics with higher geometric fidelity, and (3) it advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) GEV model that integrates roadway geometry, traffic directionality, and multilevel heterogeneity for corridor-wide yet location-sensitive crash risk estimation.
Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for proactive corridor-wide crash risk prediction, leveraging Argoverse-2 trajectory and HD map data to support both real-time interventions and long-term safety planning.
@@ 2 Near-miss detection framework
This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation. The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions. Together, these components enable the accurate forecasting of vehicle trajectories and identification of potential near misses. The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
@@ 2.1 Vehicle dynamics model
The vehicle dynamics module provides the predictive foundation of the framework. This adopts the kinematic bicycle model to represent vehicle footprint, as it is suitable for short-term near-miss analysis, where steering and acceleration dominate, and lateral slip can be neglected. The motion is represented by a four-dimensional (4D) dynamic state vector. At any time $t$, the state of vehicle $i$ is given by:
$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] \theta_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
Here, $x_{i}(t),y_{i}(t)$ denote the vehicle’s global center coordinates, $\theta_{i}(t)$ its heading angle, and $v_{i}(t)$ its speed. Over the prediction window, each vehicle is assumed to maintain constant acceleration $a_{i}$ and steering angle $\delta_{i}$, simplifying short-horizon trajectory integration. The motion then follows the standard kinematic bicycle equations:
@@
Here, $L_{i}$ is the vehicle wheelbase, influencing the turning radius and path curvature of vehicle $i$. The steering angle $\delta_{i}$ and acceleration $a_{i}$ are treated as fixed inputs over the lookahead period. Equations (2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory. To model interactions, individual states are concatenated into an eight-dimensional (8D) joint state vector, defined as:
@@
The state-space vector field in Eqn.7 gives the instantaneous rate of change of the joint state $\dot{X}(t)$. Its forward evolution is computed using the fourth-order Runge–Kutta (RK4) method, which balances accuracy and efficiency for nonlinear vehicle dynamics. The time horizon is discretized into $N$ steps of size $\Delta t$ ($t_{n}=n\Delta t$, $n=0,1,\dots,N$), and the RK4 update for state ${X}(t_{n+1})$ from ${X}(t_{n})$ is:
@@
This integration is repeated for each timestep $n\in\{0,\dots,N-1\}$, producing a sequence of predicted states $\{{X}_{n}\}$ over the horizon. These trajectories form the basis for geometric collision checks between vehicles and roadway boundaries.
@@ 2.2 Vehicle geometry representation
This transformation rotates the local corner coordinates by the heading angle and translates them into the global frame. Repeating it for all corners at each timestep produces the vehicle’s time-varying geometric footprint, which serves as input for near-miss detection. These footprints enable precise evaluation of proximity-based events, both V-V and V-I, ensuring accurate 2D-TTC estimation while reducing false positives from geometric simplifications.
@@ 2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations. At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations. A collision is defined when a vehicle’s corner lies within a proximity threshold of a corner of another vehicle. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
@@
An exact match between corners of two vehicles is highly unlikely due to floating-point arithmetic and continuous motion. Instead, a practical criterion is applied: a collision is flagged if the difference in both the x and y coordinates of any corner pair is less than or equal to a small threshold $\epsilon$:
@@
At each timestep, all 16 corner-pair combinations are evaluated. A V–V near-miss is flagged if any pair satisfies the proximity condition, and the corresponding collision time $t_{c}$ is recorded for 2D-TTC computation. This discrete-time approach enables accurate detection of spatial overlap while avoiding the oversimplifications of circular vehicle approximations.
@@ 2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near misses between vehicles and roadway infrastructure, including curbs, medians, or barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
Road boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2. At each timestep, the global positions of the four vehicle corners ${p}_{i}^{(j)}(t)$ are obtained from the predicted state. The distance between a vehicle corner $j$ and boundary point $\ell$, the distance at time $t$ is computed as:

### v4

@@ Abstract
Real-time, corridor-wide prediction of crash-occurrence risk (COR) is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to adequately account for spatial heterogeneity in traffic and roadway conditions. To address these gaps, this study develops a geometry-aware 2D-TTC near-miss extraction and integrates it with a hierarchical Bayesian structure grouped random parameters (HBSGRP–UGEV) to estimate short-term COR in urban corridors. Building on prior grouped EVT formulations while explicitly accommodating both V–V and V–I near-miss processes within a unified corridor-wide modeling framework. High-resolution trajectories from the Argoverse-2 dataset were analyzed across 28 sites on Miami’s Biscayne Boulevard to extract extreme near-miss events. The model incorporates vehicle dynamics and roadway features as covariates, with partial pooling across segments and intersections to capture corridor-wide heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% (V–V) and 3.1% (V–I). Predictive validation using ROC–AUC confirms strong accuracy (0.89 for V–V segments, 0.82 for intersections, 0.79 for V–I segments, and 0.75 for intersections). Grouped random-parameter results indicate that relative (speed, distance, and deceleration) dominate V–V near-miss risk on segments, whereas V–I segment risk is primarily associated with relative distance; at intersections, V–V risk is driven by relative (speed and distance), while V–I dynamics exhibit no statistically significant effects. These findings demonstrate the value of a geometry-aware, spatially adaptive framework for proactive corridor safety management, supporting both real-time interventions and long-term Vision Zero goals.
@@ 1 Introduction
The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches. Urban corridors are particularly vulnerable, as even a single crash can disrupt mobility, trigger cascading congestion, and result in substantial economic losses. In 2023, traffic crashes in the US claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes exceeds \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite decades of research, corridor-wide crash risk prediction remains methodologically limited.
@@
Crash risk prediction has traditionally relied on retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011). Although such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions. These drawbacks highlight the need for proactive approaches that anticipate rather than react to crash risk.
@@
To address these challenges, researchers have shifted toward proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (e.g., Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among these, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, their assumptions of constant velocity and linear paths can bias near-miss characterization, especially in turning and boundary-constrained settings. Recent two-dimensional (2D) extensions of TTC (Li et al., 2024; Anis et al., 2025b) that incorporate position, heading, yaw rate, speed, and acceleration often still rely on simplified vehicle shape approximations (e.g., circular). Although Hou et al. 2014 compared alternative vehicle-shape representations (e.g., circle, rectangle, and hybrid forms) for simplified TTC computation in simulation, a corridor-scale, high-fidelity 2D-TTC that propagates geometry-aware footprints and jointly extracts V–V and V–I near-misses from trajectories remains limited.
@@
Recent studies have integrated computed SSMs with the extreme value theory (EVT) framework to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Since EVT can extrapolate from frequent near misses or SSMs to rare crash-prone extremes, it offers a statistically principled framework for proactive safety analysis. Yet most near-miss and EVT frameworks remain limited, focusing narrowly on V–V interactions and excluding V–I near-misses such as lane edges, curbs, and barriers. This omission is critical because V–I events account for a large share of severe roadway departure crashes (Islam et al., 2021). Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable precise alignment of roadway boundaries with vehicle trajectories, creating new opportunities to model both V–V and V–I near misses with geometric fidelity.
@@
Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across the study area. Over time, nonstationary GEV formulations have been developed, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022b; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby represent variation across traffic states and locations. (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions further address data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). A growing body of EVT-based safety studies now incorporates vehicle dynamics and roadway context (Fu and Sayed, 2022b; Kumar and Mudgal, 2024; Anis et al., 2025b; Singh et al., 2025) and, in some cases, distinguishes traffic-flow directions (Kamel et al., 2024). However, direction-wise parametrization of vehicle-dynamics effects remains limited, particularly for corridor segments where interaction mechanisms can differ systematically by travel direction and access density. Moreover, most EVT applications focus on intersection settings (Fu and Sayed, 2023, 2021c, 2022b; Ghoul et al., 2023a; Kumar and Mudgal, 2024), only a small subset extends to corridor or network-scale inference (Ghoul et al., 2023b; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Anis et al., 2025b; Singh et al., 2025).
To address these gaps, this paper makes two contributions. First, it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions by implementing a geometry-aware 2D-TTC formulation that forward-projects rectangular vehicle footprints and uses HD-map boundaries to define V–I exposure, enabling joint extraction of V–V and V–I extremes from AV trajectories. Second, it adapts and extends a hierarchical Bayesian structure grouped random-parameters (HBSGRP) UGEV framework for corridor-scale modeling, integrating fixed roadway and time-varying vehicle-dynamics covariates within an interaction-window BM framework to capture multilevel heterogeneity across intersections and directional segments. Together, these contributions establish an infrastructure-aware and spatially adaptive framework for proactive corridor-wide crash-risk estimation that links high-fidelity near-miss geometry with hierarchical EVT inference, supporting both real-time intervention and longer-term safety planning.
@@ 2 Near-miss detection framework
This section presents the computational framework for estimating a geometry-aware, high-fidelity 2D-TTC to support real-time COR inference. The proposed workflow follows projection-based TTC concepts that forward-simulate oriented rectangular footprints, but is implemented directly on high-frequency AV trajectories and HD-map boundary primitives. The framework integrates a continuous-time kinematic bicycle model, explicit vehicle footprint geometry, and a spatial proximity algorithm to capture both V–V and V–I interactions. Vehicles are represented as oriented bounding boxes, and V–I exposure is defined using the curb/median boundaries of the vehicle’s current carriageway extracted from the HD map, enabling consistent joint detection of near misses along the corridor. Together, these components provide a deterministic, high-resolution basis for extracting extreme near-miss severity used in the subsequent EVT-based COR model.
@@ 2.1 Vehicle dynamics model
The vehicle dynamics module provides the predictive foundation of the framework. Which adopts the kinematic bicycle model for short-horizon interaction forecasting, where steering and acceleration dominate, and lateral tire slip can be neglected. The motion is represented by a four-dimensional (4D) dynamic state vector. At any time $t$, the state of vehicle $i$ is given by Eq. (1):
$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
Here, $x_{i}(t),y_{i}(t)$ denote the vehicle’s global center coordinates, $\theta_{i}(t)$ its heading angle, and $v_{i}(t)$ its speed. Over the prediction window, each vehicle is assumed to maintain constant acceleration $a_{i}$ and steering angle $\delta_{i}$, simplifying short-horizon trajectory integration. The motion then follows the standard kinematic bicycle Eqs. (2-5):
@@
Here, $L_{i}$ is the vehicle wheelbase, influencing the turning radius and path curvature of vehicle $i$. The steering angle $\delta_{i}$ and acceleration $a_{i}$ are treated as fixed inputs over the lookahead period. Eqs. (2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory. To model interactions, individual states are concatenated into an eight-dimensional (8D) joint state vector, defined as:
@@
The state-space vector field in Eq. (7) gives the instantaneous rate of change of the joint state $\dot{X}(t)$. Its forward evolution is computed using the fourth-order Runge–Kutta (RK4) method, which balances accuracy and efficiency for nonlinear vehicle dynamics. The time horizon is discretized into $N$ steps of size $\Delta t$ ($t_{n}=n\Delta t$, $n=0,1,\dots,N$), and the RK4 update for state ${X}(t_{n+1})$ from ${X}(t_{n})$ is:
@@
This integration is repeated for each timestep $n\in\{0,\dots,N-1\}$, producing a sequence of predicted states $\{{X}_{n}\}$ over the horizon. These predicted states are then converted to oriented footprints for geometric proximity checks against other vehicles and HD-map boundaries.
@@ 2.2 Vehicle geometry representation
This transformation rotates the local corner coordinates by the heading angle and translates them into the global frame. Repeating it for all corners at each timestep produces the vehicle’s time-varying geometric footprint, which serves as input for near-miss detection. This footprint construction is applied at every RK4 step, yielding a time-varying oriented bounding box suitable for both V–V and V–I proximity evaluation.
@@ 2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and predicted states and orientations determine their interactions. At each RK4 integration timestep, the global positions of all four corners are computed using the kinematic and rotational transformations. A V–V near-miss is detected when the predicted footprints exhibit near-contact under a numerical tolerance $\epsilon$, implemented through corner-based proximity checks across all corner pairs. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
@@
An exact match between the corners of two vehicles is highly improbable, given floating-point arithmetic and continuous motion. Instead, a practical criterion is applied: a collision is flagged if the difference in both the x and y coordinates of any corner pair is less than or equal to a small threshold $\epsilon$:
@@
At each timestep, all 16 corner-pair combinations are evaluated. A V–V near-miss is flagged if any pair satisfies the proximity condition, and the corresponding collision time $t_{c}$ is recorded for 2D-TTC computation. This discrete-time footprint-based check aligns with high-frequency trajectory sampling and avoids the geometric bias introduced by circular approximations in oblique and turning interactions.
@@ 2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure, including curbs, medians, and barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
In this study, the scope is defined as the curb and median boundaries of the vehicle’s current carriageway as encoded in the HD map; opposing-direction curbs, minor side-street sidewalks, and internal lane markings are beyond the scope. To do so, curb and boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2. At each timestep, the global positions of the four vehicle corners ${p}_{i}^{(j)}(t)$ are obtained from the predicted state. The distance between a vehicle corner $j$ and boundary point $\ell$, the distance at time $t$ is computed as: